State whether the following statements are true or false. Justify yourself with the help of a short proof or a counter example. (2) The solution of PDE dz/dx+dz/dy=z^2 is z=-[y+f(x-y) ].
1
Expert's answer
2020-03-16T13:32:27-0400

It is false.

If we take "f(u)=0", then we obtain "z=-y".

Then "\frac{dz}{dx}=0", "\frac{dz}{dy}=-1" and "z^2=y^2".

Functions "0+(-1)=-1" and "y^2" are not equal.